OpenStudy (anonymous):

solve: ln(x^2-x-6)=7+ln(x+2)

OpenStudy (dumbcow):

x^2 -x-6 = (x-3)(x+2) ln[(x-3)(x+2)] = ln(x-3) + ln(x+2) --> ln(x-3) = 7 --> x-3 = e^7 -->x = e^7 +3

OpenStudy (anonymous):

ty so much